OpenStudy (anonymous):

Roberto is going to visit his family in Mexico. He is bringing $500 US dollars. On the way over, he can exchange his US Dollars for Pesos. The exchange rate is $0.07 USD = 1 Peso. How much money will he have to spend in Mexico, according to their monetary system? A. 35 Pesos B. 700 Pesos C. $0.0004 Pesos D. $7142.86 Pesos I got B as my answer, would that be correct?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

0.07 dollars/1 peso = 500 dollars/x pesos 0.07/1 = 500/x 0.07x = 500*1 0.07x = 500 x = 500/0.07 x = ???

OpenStudy (anonymous):

its D

jimthompson5910 (jim_thompson5910):

better

OpenStudy (anonymous):

so would that be my answer than?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

acyua;;y im sorry those are the wrong multiple choices

jimthompson5910 (jim_thompson5910):

huh? how so

OpenStudy (anonymous):

A. $35 USD/peso B. $700 USD/ pesp C. 0.07 USD/ peso D. 0.0004 Pesos Because I tried coping and pasting the question so I wouldnt have to type it all

jimthompson5910 (jim_thompson5910):

ah ok

OpenStudy (anonymous):

so I got B for this one

jimthompson5910 (jim_thompson5910):

same problem?

OpenStudy (anonymous):

yep

jimthompson5910 (jim_thompson5910):

0.07 dollars = 1 peso how much is $500 worth?

OpenStudy (anonymous):

no lol sory

jimthompson5910 (jim_thompson5910):

i was gonna say...

jimthompson5910 (jim_thompson5910):

lol

OpenStudy (anonymous):

ill rewrite the worded problem lol

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

Roberto is going to visit his family in Mexico. He is bringing $500 US dollars. On the way over, he can exchange his US Dollars for Pesos and got $7142.86. what is the exchange rate USD to pesos?

jimthompson5910 (jim_thompson5910):

hmm it's like they flipped it around

OpenStudy (anonymous):

ya lol

jimthompson5910 (jim_thompson5910):

wait...did you mean "he can exchange his US Dollars for Pesos and got 7142.86 pesos" ?

OpenStudy (anonymous):

yes but without the peso at the end

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

you just need to first calculate 500/7142.86 to get 500/7142.86 = ?? and you'll need to round as well

OpenStudy (anonymous):

okay so our answer would be C then correct?

jimthompson5910 (jim_thompson5910):

yes, C. 0.07 USD/ peso

jimthompson5910 (jim_thompson5910):

which was given in the last problem

OpenStudy (anonymous):

yep! thanks

jimthompson5910 (jim_thompson5910):

yw